OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line which passes through the points (1, 6) and (2, 9).

OpenStudy (kamibug):

We will put it in point-slope form and then convert it. First find the slope using the slope formula. \[m = \frac{ y_2 - y_1 }{ x_2 - x_1 }\] m = 9 - 6 / 2 - 1 m = 3 / 1 m = 3 Plug the slope and one of the points into point-slope form. y - 9 = 3(x - 2) Now we'll convert this equation to slope-intercept form. Distribute. y - 9 = 3x - 6 y = 3x + 3 <-- this is your answer. :)

OpenStudy (anonymous):

thank you